# Decoding Neural Emotion Patterns through Natural Language Processing Embeddings

### v1

@@ 3.1 Experiment 1: Healthy versus Depressed Subjects
Amygdala Left 0.1000 0.1000 10.0000 1.0000 No
Amygdala Right 0.1000 0.1000 20.0000 1.0000 No
Anterior Cingulate Right 0.1000 0.1000 6.0000 1.0000 No
Insula Left 0.1000 0.0781 37.5000 0.0041 Yes
Insula Right 0.0950 0.1000 12.0000 0.5002 No
Orbitofrontal Left 0.1000 0.1000 4.5000 1.0000 No
Hippocampus Left 0.1000 0.1000 6.0000 1.0000 No
Temporal Pole Left 0.1000 0.1000 9.0000 1.0000 No
Temporal Pole Right 0.0950 0.1000 10.0000 0.4237 No
Superior Temporal Left 0.1000 0.1000 3.0000 1.0000 No
Superior Temporal Right 0.1000 0.1000 2.0000 1.0000 No
Caudate Right 0.0917 0.1000 3.0000 0.5050 No
Putamen Left 0.1000 0.1000 6.0000 1.0000 No
Putamen Right 0.1000 0.1000 10.0000 1.0000 No
Nucleus Accumbens Left 0.1000 0.1000 2.0000 1.0000 No
Raphe Nuclei 0.1000 0.0750 16.0000 0.0131 Yes
Posterior Cingulate 0.1000 0.1000 3.0000 1.0000 No
@@ 3.2 Experiment 2: Multiple Emotional States
The emotion intensity analysis results revealed a hierarchy of affective experiences, with love emerging as the most intense emotion (0.709), followed by joy (0.593) and relief (0.560). Negative emotions like sadness (0.486), fear (0.412), and anger (0.390) occupy middle-intensity positions. This intensity hierarchy suggests that basic positive emotions tend to be experienced more intensely than negative ones, with love showing remarkably high activation (Figure 6). The data also indicates that socially-oriented emotions (love, gratitude, curiosity) and approach-motivated states (joy, excitement) generate stronger neural responses than avoidance-motivated emotions (fear, disgust) or complex cognitive emotions requiring more nuanced processing (Figure 7).
Figure 6: Emotion intensity hierarchy from high activation count (left) to lower activation count (right).
@@ 3.3 Experiment 3: Human versus LLM Chat bot
Comparing the results of human textual conversation analysis with the responses generated by an LLM chat bot (Figure 8), the analysis revealed distinct activation pattern differences between human and LLM conversational profiles. Human texts demonstrated relatively balanced engagement across regions, with moderate amygdala activation suggesting appropriate emotional regulation [64, 65], balanced prefrontal cortex activity indicating cognitive control [66, 67], and measured caudate activation reflecting normal reward processing [68, 69]. Additionally, humans showed moderate orbito-frontal activation consistent with healthy social cognition [70, 71].
In contrast, the LLM-generated responses exhibited markedly different patterns, including heightened activation in several regions such as the superior temporal areas, caudate, and orbito-frontal regions, alongside notably reduced activation in areas like the putamen and ventral tegmental area. These patterns suggest different underlying computational processes between human language generation and LLM text production [10, 11].
Figure 8: Human vs. LLM chat bot emotion activation count comparison.
Our proposed approach shows promise in distinguishing human-authored text from LLM-generated content, supporting recent studies [31, 30, 28, 29] that have demonstrated the potential of computational approaches in analyzing text to predict and classify various characteristics. These results suggest that natural language embeddings may encode information beyond surface-level semantics that correlates with different processing patterns.
Figure 9 shows a 3D cortical rendering in MNI space, with lateral (left) and medial (right) views indicating the magnitude of differential activation between human subjects and LLM chat bot responses (Table 3). The visualization represents computationally derived activation patterns, where embeddings originally in 1536-dimensional space were reduced to three principal components using PCA and spatially projected onto the cortical surface. Red shading indicates the magnitude of differential activity captured by the model, with distinct patterns evident between humans and the LLM across multiple cortical regions.
Figure 9: 3D rendering of the caudal (left) and parietal (right) regions with emotion intensity differences (Table 3) between human subjects and an LLM chat bot shown in red.
@@
Amygdala Left 0.3027 0.2563 85770.5 0.0154 Yes
Amygdala Right 0.3114 0.2047 81751.0 0.0001 Yes
Anterior Cingulate Left 0.3385 0.5938 43113.0 0.0001 Yes
Anterior Cingulate Right 0.2588 0.1204 182507.0 0.0001 Yes
Insula Left 0.1825 0.2668 26966.5 0.0001 Yes
Insula Right 0.3236 0.1371 131640.5 0.0001 Yes
Orbitofrontal Left 0.4074 0.2805 99956.0 0.0001 Yes
Orbitofrontal Right 0.1761 0.2825 33757.0 0.0001 Yes
Hippocampus Left 0.3129 0.3085 101184.0 0.0004 Yes
Hippocampus Right 0.1855 0.3724 65936.5 0.0001 Yes
Prefrontal Cortex Left 0.1823 0.6846 12674.5 0.0001 Yes
Prefrontal Cortex Right 0.1781 0.4598 16839.0 0.0001 Yes
Temporal Pole Left 0.2686 0.1453 122482.5 0.0001 Yes
Temporal Pole Right 0.2785 0.2922 47292.5 0.0001 Yes
Superior Temporal Left 0.3130 0.1240 86434.5 0.0001 Yes
Superior Temporal Right 0.3648 0.2493 89242.5 0.0001 Yes
Caudate Left 0.2425 0.1508 67663.0 0.0001 Yes
Caudate Right 0.4373 0.2392 16469.5 0.0001 Yes
Putamen Left 0.2063 0.4327 68661.0 0.0001 Yes
Putamen Right 0.2187 0.1139 103474.0 0.0001 Yes
Nucleus Accumbens Left 0.1867 0.1351 190880.0 0.0001 Yes
Nucleus Accumbens Right 0.2072 0.1886 48552.5 0.0004 Yes
Hypothalamus 0.2885 0.2684 25680.5 0.6085 No
Periaqueductal Gray 0.2471 0.2699 53214.0 0.0001 Yes
Ventral Tegmental Area 0.3154 0.3274 65572.5 0.00001 Yes
@@ 4 Study Limitations
Several important limitations must be acknowledged. Firstly, the mapping from text embeddings to brain regions represents a computational model rather than direct measurement of neural activity. Secondly, further empirical validation through direct comparison with neuroimaging data, such as fMRI, would be necessary to establish the neurobiological validity of these mappings. Although recent advances have shown that brain signals from fMRI and EEG can be decoded into coherent text [72, 73, 74], the inverse problem of predicting potentially engaged brain regions during language processing based on text input remains largely unexplored and requires careful validation. Finally, the predefined regional coordinates, while based on neuroimaging research, represent population averages that may not accurately reflect individual neuro-anatomy.
@@ 5 Conclusion
The proposed approach represents a significant advancement in our ability to study emotional processing through computational methods. By leveraging state-of-the-art natural language processing techniques and established neuro-anatomical knowledge, the proposed approach offers a scalable, accessible alternative to traditional neuroimaging methods. Our proposed approach’s ability to process natural language and map emotional content onto anatomically defined brain regions opens new possibilities for understanding individual differences in emotional processing, monitoring mental health at scale, and developing personalized interventions.

### v2

@@ 3.1 Experiment 1: Healthy versus Depressed Subjects
Amygdala Left 0.1000 0.1000 8.0000 1.000 N
Amygdala Right 0.1000 0.1000 30.0000 1.000 N
Anterior Cingulate Left 0.1000 0.1000 N
Anterior Cingulate Right 0.1000 0.1000 6.0000 1.000 N
Insula Left 0.1000 0.0781 52.5000 0.001 Y
Insula Right 0.0964 0.1000 9.0000 0.662 N
Orbitofrontal Left 0.1000 0.1000 4.5000 1.000 N
Orbitofrontal Right 0.1000 0.1000 3.0000 1.000 N
Hippocampus Left 0.1000 0.1000 6.0000 1.000 N
Hippocampus Right 0.0750 0.1000 N
Prefrontal Cortex Left 0.1000 0.1000 N
Temporal Pole Left 0.1000 0.1000 6.0000 1.000 N
Temporal Pole Right 0.0917 0.1000 5.0000 0.301 N
Superior Temporal Left 0.1000 0.1000 2.0000 1.000 N
Superior Temporal Right 0.1000 0.1000 3.0000 1.000 N
Caudate Left 0.1000 0.1000 N
Caudate Right 0.0917 0.1000 3.0000 0.504 N
Putamen Left 0.0938 0.1000 4.5000 0.563 N
Putamen Right 0.1000 0.1000 5.0000 1.000 N
Nucleus Accumbens Left 0.1000 0.1000 3.0000 1.000 N
Nucleus Accumbens Right 0.1000 0.1000 N
Hypothalamus 0.1000 0.0750 N
Periaqueductal Gray 0.1000 0.1000 10.0000 1.000 N
Raphe Nuclei 0.1000 0.0750 16.0000 0.013 Y
Locus Coeruleus 0.1000 0.1000 N
Posterior Cingulate 0.1000 0.1000 N
Medial Prefrontal Cortex 0.1000 0.1000 4.0000 1.000 N
The broader analysis by brain system categories (Table 2) revealed systematic differences in activation patterns. Cortical regions showed the most substantial difference, with healthy subjects displaying 40 total activations compared to 13 in depressed subjects (a 67% reduction). Subcortical regions showed healthy subjects with 32 activations versus 14 in depressed subjects (a 56% reduction). The limbic system demonstrated the smallest absolute difference, with 23 activations in healthy subjects compared to 12 in depressed subjects, though this still represents a 48% reduction.
Figure 5 presents 3D cortical surface renderings in MNI space, comparing healthy (left) and depressed (right) groups. In the lateral views (top), healthy subjects exhibit robust bilateral activation across the lateral occipital cortices, whereas depressed subjects show reduced intensity and spatial extent in the same regions. In the dorsal views (bottom), activation in the healthy group spans broadly across the posterior occipital cortices, while depressed participants again display markedly diminished engagement. This pattern points to altered visual processing network function in depression, consistent with prior reports of occipital cortical abnormalities, including disrupted dynamics at rest and altered connectivity with emotion-regulation systems [65, 66, 67].
Figure 5: 3D rendering of emotion intensity activation differences (Table 3) showing lateral (top) and dorsal (bottom) views between healthy (left) and depressed (right) subjects shown in red.
The multi-trial validation analysis (discussed in section 2.8) revealed a systematic difference in emotional response patterns between groups (Table 4). Healthy participants demonstrated variable silhouette scores ranging from 0.20 - 0.39 across strategies, indicating heterogeneous emotional expression patterns. In contrast, depressed participants consistently exhibited higher silhouette scores (0.27 - 0.89), suggesting homogeneous, constrained emotional response patterns.
The highlighted regions in Table 4 were selected based on three convergent criteria: i) highest clustering quality ratios, ii) established roles in depression pathophysiology, and iii) representation of distinct functional brain systems. This approach ensured both statistical robustness and neurobiological interpretability.
Table 4: Multi-Trial Analysis Results: Original vs. Multi-Trial
Region Original (Imbalanced) Balanced Multi-Trial Clustering Quality
H Mean D Mean p-value Mean p Sig Trials Effect Size H Silhouette D Silhouette Quality Ratio
(n=97) (n=37) (MWU) (15 trials) (%) (Cohen’s d) Range Range (D/H)
OVERALL SUMMARY
Global Intensity 1.997 1.984 0.122 0.297 6.7% 0.89±0.31 0.20-0.39 0.27-0.89 2.23x
LIMBIC SYSTEM
Amygdala Left 0.100 0.100 1.000 0.892 0% 0.12±0.08 0.18-0.34 0.25-0.82 2.28x
Amygdala Right 0.100 0.100 1.000 0.847 0% 0.15±0.11 0.19-0.35 0.26-0.84 2.31x
Anterior Cingulate Left 0.100 0.100 0.923 0% 0.08±0.05 0.17-0.33 0.24-0.79 2.18x
Anterior Cingulate Right 0.100 0.100 1.000 0.856 0% 0.14±0.09 0.18-0.34 0.25-0.81 2.25x
Hippocampus Left 0.100 0.100 1.000 0.789 6.7% 0.18±0.13 0.20-0.36 0.27-0.85 2.36x
Hippocampus Right 0.075 0.100 0.734 13.3% 0.22±0.15 0.19-0.35 0.28-0.86 2.42x
CORTICAL REGIONS
Insula Left 0.100 0.078 0.001 0.245 26.7% 0.45±0.18 0.22-0.41 0.31-0.92 2.68x
Insula Right 0.096 0.100 0.662 0.678 6.7% 0.19±0.12 0.21-0.38 0.29-0.87 2.44x
Orbitofrontal Left 0.100 0.100 1.000 0.912 0% 0.10±0.07 0.18-0.33 0.24-0.78 2.15x
Orbitofrontal Right 0.100 0.100 1.000 0.889 0% 0.11±0.08 0.17-0.32 0.23-0.77 2.12x
Prefrontal Cortex Left 0.100 0.100 0.834 0% 0.16±0.10 0.19-0.35 0.26-0.83 2.33x
Medial Prefrontal Cortex 0.100 0.100 1.000 0.798 6.7% 0.17±0.12 0.20-0.37 0.28-0.88 2.39x
Temporal Pole Left 0.100 0.100 1.000 0.867 0% 0.13±0.09 0.18-0.34 0.25-0.81 2.22x
Temporal Pole Right 0.092 0.100 0.301 0.645 6.7% 0.20±0.13 0.20-0.37 0.28-0.86 2.41x
Superior Temporal Left 0.100 0.100 1.000 0.889 0% 0.11±0.08 0.17-0.33 0.24-0.79 2.18x
Superior Temporal Right 0.100 0.100 1.000 0.901 0% 0.10±0.07 0.17-0.32 0.23-0.77 2.14x
SUBCORTICAL REGIONS
Caudate Left 0.100 0.100 0.923 0% 0.09±0.06 0.17-0.32 0.23-0.76 2.11x
Caudate Right 0.092 0.100 0.504 0.698 6.7% 0.18±0.12 0.19-0.36 0.27-0.84 2.35x
Putamen Left 0.094 0.100 0.563 0.712 6.7% 0.17±0.11 0.19-0.35 0.26-0.83 2.32x
Putamen Right 0.100 0.100 1.000 0.845 0% 0.15±0.10 0.18-0.34 0.25-0.82 2.28x
Nucleus Accumbens Left 0.100 0.100 1.000 0.878 0% 0.12±0.08 0.18-0.33 0.24-0.80 2.20x
Nucleus Accumbens Right 0.100 0.100 0.912 0% 0.10±0.07 0.17-0.32 0.23-0.77 2.15x
Hypothalamus 0.100 0.075 0.298 20.0% 0.43±0.17 0.22-0.40 0.31-0.91 2.65x
Periaqueductal Gray 0.100 0.100 1.000 0.834 0% 0.16±0.10 0.19-0.35 0.26-0.83 2.33x
Raphe Nuclei 0.100 0.075 0.013 0.189 33.3% 0.52±0.21 0.23-0.42 0.32-0.95 2.71x
Ventral Tegmental Area 0.089 0.100 0.445 0.567 13.3% 0.24±0.14 0.21-0.39 0.30-0.89 2.48x
Locus Coeruleus 0.100 0.100 0.923 0% 0.09±0.06 0.17-0.32 0.23-0.76 2.11x
Posterior Cingulate 0.100 0.100 0.867 0% 0.13±0.09 0.18-0.34 0.25-0.81 2.22x
The clustering quality ratio (depressed vs. healthy silhouette scores) revealed that depressed participants showed 2.2 - 2.3× more homogeneous clustering patterns across all balancing strategies, indicating a robust, sample-size independent difference in emotional pattern diversity. The analysis further revealed a previously unrecognized difference between healthy and depressed populations: depressed individuals demonstrate emotional pattern rigidity, a constraint in the diversity and flexibility of emotional responses [68]. This finding aligns with emerging theories of depression emphasizing cognitive and behavioral inflexibility [69, 70], and extends these concepts to emotional processing patterns derived from natural language. The clustering pattern differences are also consistent with neuroimaging findings showing reduced network flexibility in depression [71, 72], while the mapping of constrained emotional patterns to brain regions aligns with evidence of altered connectivity and reduced neural network switching in depressed individuals [73, 74].
The high silhouette scores in the depressed group suggest stereotyped, predictable emotional expressions, while the variable clustering in healthy participants indicates adaptive emotional flexibility: the ability to express emotions across a broader range of patterns depending on context [75]. This distinction has important clinical implications, suggesting that therapeutic interventions might benefit from targeting emotional range expansion rather than intensity modification alone. This finding suggests that assessment tools should evaluate emotional diversity and flexibility rather than focusing solely on intensity or valence measures. Finally, assessment tools such as machine learning classification models that rely on text for health screening, should account for emotional diversity and flexibility, rather than focusing solely on intensity or valence measures.
@@ 3.2 Experiment 2: Multiple Emotional States
The second experiment was performed on the GoEmotions dataset [38], which includes 58,000 Reddit [39] comments manually labeled into 27 emotion categories (or neutral). The emotion intensity analysis using our method revealed a hierarchy of affective experiences, with love emerging as the most intense emotion (0.709), followed by joy (0.593) and relief (0.560). Negative emotions like sadness (0.486), fear (0.412), and anger (0.390) occupy middle-intensity positions. This intensity hierarchy suggests that basic positive emotions tend to be experienced more intensely than negative ones, with love showing remarkably high activation (Figure 6). The data also indicates that socially-oriented emotions (love, joy, relief) and approach-motivated states (excitement) generate stronger neural responses than avoidance-motivated emotions (fear, disgust) or complex cognitive emotions requiring more nuanced processing (Figure 7).
Figure 6: Scaled motion intensity hierarchy from high activations (left) to lower activations (right).
@@ 3.3 Experiment 3: Human versus LLM Chatbot
The third experiment was performed on the Schema-Guided Dialogue dataset [40], which represents nearly half a million sentences comprised of human and LLM chatbot interactions. Comparing human conversational texts with LLM-generated responses revealed systematic divergences in neural activation profiles across limbic, cortical, and brainstem regions. Figure 8 shows a 3D cortical rendering in MNI space, with lateral (top) and dorsal (bottom) views indicating the magnitude of differential activation between human subjects (left) and LLM chatbot (right) responses. The visualization represents computationally derived activation patterns, where embeddings originally in 1536-dimensional space were reduced to three principal components using PCA and spatially projected onto the cortical surface. Red shading indicates the magnitude of differential activity captured by the model, with distinct patterns evident between humans and the LLM across multiple cortical regions.
Figure 8: 3D rendering of activation differences (Table 5) showing lateral (top) and dorsal (bottom) views between human subjects (left) and an LLM chatbot (right).
@@
Amygdala Left 0.3240 0.1695 145278.0000 0.000 Y
Amygdala Right 0.3447 0.2667 26617.0000 0.008 Y
Anterior Cingulate Left 0.3004 0.7196 15865.5000 0.000 Y
Anterior Cingulate Right 0.3071 0.4716 32510.5000 0.000 Y
Insula Left 0.2633 0.3037 44791.5000 0.143 N
Insula Right 0.2171 0.2671 0.000 Y
Orbitofrontal Left 0.1841 0.3101 110574.5000 0.000 Y
Orbitofrontal Right 0.3181 0.1120 86870.0000 0.000 Y
Hippocampus Left 0.2636 0.2100 103236.5000 0.032 Y
Hippocampus Right 0.2299 0.6703 19439.0000 0.000 Y
Prefrontal Cortex Left 0.1681 0.2687 40032.5000 0.000 Y
Prefrontal Cortex Right 0.2212 0.2906 37838.5000 0.000 Y
Temporal Pole Left 0.4250 0.3322 58079.5000 0.000 Y
Temporal Pole Right 0.3700 0.1330 61866.0000 0.000 Y
Superior Temporal Left 0.1782 0.1404 125718.5000 0.000 Y
Superior Temporal Right 0.2805 0.2603 13359.0000 0.979 N
Caudate Left 0.1911 0.1183 76483.5000 0.000 Y
Caudate Right 0.2189 0.1217 86438.5000 0.000 Y
Putamen Left 0.3266 0.4805 30059.0000 0.000 Y
Putamen Right 0.3975 0.2521 80306.0000 0.000 Y
Nucleus Accumbens Left 0.1976 0.2876 20190.0000 0.000 Y
Nucleus Accumbens Right 0.2298 0.2579 0.000 Y
Hypothalamus 0.2812 0.1316 40237.5000 0.000 Y
Periaqueductal Gray 0.2041 0.4805 30230.0000 0.000 Y
Ventral Tegmental Area 0.3640 0.1611 72183.0000 0.000 Y
Raphe Nuclei 0.3639 0.2364 10554.0000 0.000 Y
Locus Coeruleus 0.1849 0.2493 0.000 Y
Posterior Cingulate 0.2231 0.1606 39274.0000 0.000 Y
Medial Prefrontal Cortex 0.2774 0.2795 0.000 Y
Humans demonstrated stronger recruitment of emotion-related regions, including bilateral amygdalae [79, 80], as well as memory-related structures such as the left hippocampus, consistent with reliance on autobiographical retrieval during dialogue [81]. The human text also showed greater engagement of reward and arousal-related circuits, including the ventral tegmental area, raphe nuclei, and locus coeruleus, reflecting dopaminergic and serotonergic modulation of motivation and adaptive arousal [82, 83]. Increased activity in medial prefrontal cortex, posterior cingulate, and temporal poles further supports integration of self-referential and affective context into human conversation [84, 85].
In contrast, LLM-generated responses showed heightened anterior cingulate activity bilaterally, aligning with its role in monitoring and conflict regulation [86, 49]. LLMs also engaged the right hippocampus more strongly, suggesting an episodic-associative rather than autobiographical memory profile [81]. Cortical valuation and decision-making appeared lateralized, with left orbitofrontal cortex stronger in LLMs, while right orbitofrontal cortex was greater in humans [87, 88]. Similarly, the putamen showed a split pattern (left stronger in LLMs, right stronger in humans). The superior temporal gyri was not stronger in LLMs. Instead, the left was more active in humans, consistent with its role in speech and semantic processing [89, 90]. Finally, LLMs exhibited modestly elevated right insula activation, possibly reflecting altered interoceptive-like signal representations.
Together, these findings indicate that humans preferentially engage limbic brainstem networks integrating affect, memory, and motivation into language use, whereas LLMs display a bias toward cingulate, orbitofrontal, and striatal pathways associated with conflict monitoring and associative sequencing. This dissociation is consistent with recent work contrasting artificial and biological language networks [10, 11].
Our proposed approach therefore shows promise in distinguishing human-authored text from LLM-generated content, supporting recent studies [32, 31, 29, 30] that have demonstrated the potential of computational approaches in analyzing text to predict and classify various characteristics. These results suggest that natural language embeddings may encode information beyond surface-level semantics that correlates with different processing patterns.
@@ 4 Study Limitations
Several important limitations must be acknowledged. Firstly, the mapping from text embeddings to brain regions represents a computational model rather than direct measurement of neural activity. Recent advances in computational neuroscience have demonstrated the feasibility of embedding-to-brain mappings: for example, Goldstein et al. [26] established that linear encoding models can effectively predict neural activity from high-dimensional text embeddings, achieving correlations up to 0.50 with actual neural signals during natural conversations. While this provides methodological support for computational approaches, our study does not directly validate these mappings against neuroimaging data.
Secondly, although brain signals from fMRI and EEG can be decoded into coherent text [91, 92, 93], the inverse problem of predicting potentially engaged brain regions during language processing based on text input remains largely unexplored and requires careful validation. Goldstein et al. [26] highlighted both the promise and the challenge of this approach, demonstrating systematic individual and regional differences in how embedding-based models align with neural activity. These findings underscore the importance of extending our population-level predictions to individual-level validation across different modalities.
Finally, the predefined regional coordinates used in our study, while grounded in established neuroimaging research, represent population averages that may not accurately capture individual neuro-anatomy or the mixed selectivity patterns revealed in prior work [26]. Future studies should address these limitations by combining embedding-based predictions with direct neuroimaging measures and by accounting for individual variability in brain-language mappings, particularly in atypical or clinical populations.
@@ 5 Conclusion
In this study, we developed a novel decoding scheme that maps neural emotion patterns to brain regions through LLM embeddings. The proposed approach represents a significant advancement in our ability to study emotional processing through computational methods. By leveraging state-of-the-art natural language processing techniques and established neuro-anatomical knowledge, the proposed approach offers a scalable, accessible alternative to traditional neuroimaging methods. Our proposed approach’s ability to process natural language and map emotional content onto anatomically defined brain regions opens new possibilities for understanding individual differences in emotional processing, monitoring mental health at scale, and developing personalized interventions.